1. What is the acceleration cause by a 5-N force on a 6-kg object?

2. What force is needed to accelerate a 98-N object 4 m/s^2?

3. What force ia needed to accelerafe a 65-kg object from rest to 30 m/s in 5s?
Answer #1

1) a= F/m= 5/ 6 = 0.833m/s^2

2) mg = 98

g= 98/9.8= 10Kg

F= ma= 10 x 4= 40N

3) a= v-u/t

a= 30 - 0/ 5= 6m/s^2

F= ma= 65 x 6= 390N